Rewrite the middle term as the sum of two terms and then factor by grouping.
x^2 – 2x – 63

Answer by praseenakos@yahoo.com(507)

Question:

Rewrite the middle term as the sum of two terms and then factor by grouping.
x^2 – 2x – 63

Answer:

Here find out two numbers, whose sum is -2{that is coefficient of 'x' term and whose product is -63(that is constant term)

Such two numbers are -9 and +7
(
Now rewrite the given expression as follows...

x^2 - 9x + 7x - 63 = 0


Now group the terms...




(x^2 - 9x) + (7x - 63) = 0



Now take take out the common term from both the groups...



==> x( x-9) + 7(x-9) = 0


Now (x-9) is common in both the terms...



==> (x-9) (x-7)



So the factors of given expression are (x-9) and (x-7).



Hope you found the explanation useful....



Regards...



Praseena.